# DICE: Staleness-Centric Optimizations for Parallel Diffusion MoE Inference

### v2

@@ Abstract
Mixture-of-Experts-based (MoE-based) diffusion models demonstrate remarkable scalability in high-fidelity image generation, yet their reliance on expert parallelism introduces critical communication bottlenecks. State-of-the-art methods alleviate such overhead in parallel diffusion inference through computation-communication overlapping, termed displaced parallelism. However, we identify that these techniques induce severe staleness-the usage of outdated activations from previous timesteps that significantly degrades quality, especially in expert-parallel scenarios. We tackle this fundamental tension and propose DICE, a staleness-centric optimization framework with a three-fold approach: (1) Interweaved Parallelism introduces staggered pipelines, effectively halving step-level staleness for free; (2) Selective Synchronization operates at layer-level and protects layers vulnerable from staled activations; and (3) Conditional Communication, a token-level, training-free method that dynamically adjusts communication frequency based on token importance. Together, these strategies effectively reduce staleness, achieving 1.26×1.26$\times$1.26 × speedup with minimal quality degradation. Empirical results establish DICE as an effective and scalable solution. Our code is publicly available at https://anonymous.4open.science/r/DICE-FF04
Figure 1: We introduce DICE, a framework designed to enhance scalability and efficiency in MoE-based diffusion models. Existing asynchronous methods accelerate parallel inference but suffer from outdated activations (staleness). DICE tackles this issue with step-, layer-, and token-level optimizations, reducing staleness and achieving up to 1.26×1.26$\times$1.26 × speedup with minimal quality loss. Left: Architectural comparison; Right: Visual quality results.
**footnotetext: Equal contribution
@@ 1 Introduction
Figure 2: Execution flows of (a) synchronous expert parallelism (no staleness); (b) displaced variant of expert parallelism(two-step staleness, displaced parallelism was proposed in DistriFusion) [20] and (c) our interweaved parallelism (one-step staleness). Each color denotes a different layer, showing how interweaving reduces staleness by staggering operations within the same step. The blue path highlights layer $N$, which spans two more steps in displaced parallelism but only one in interweaved. FID and speedup refer to interweaved parallelism alone (Sec. 5.5).
Diffusion models [12, 3] have revolutionized the field of generative modeling, enabling the creation of high-fidelity images [14, 34] and videos [13, 23], rivaling human creativity in quality and detail. Among the architectures, the Diffusion Transformer (DiT) [28, 35] with Mixture-of-Experts (MoE) [32] stands out due to its effectiveness in scalability. MoE enables large diffusion models to scale up with sub-linear cost increases. It divides the model into multiple experts, with only a subset activated for each input. This selective activation reduces computation, achieving high performance at a lower cost. As demonstrated by DiT-MoE [5, 38], MoE-based diffusion models have been scaled to 16 billion parameters with superior quality.
Scaling MoE-based diffusion models necessitates expert parallelism [16, 29] to handle their extensive memory footprint; however, it also introduces significant communication bottlenecks. By distributing expert weights across multiple GPUs, expert parallelism substantially reduces memory usage but incurs significant communication overhead [9, 17, 39, 18]. As outlined in Figure 2(a), two all-to-all communications (dispatch + combine) are required per layer. This bottleneck presents a major challenge for scaling diffusion models.
To address communication challenges in parallel inference, state-of-the-art methods employ Displaced Parallelism [20]. In standard synchronous execution, communication blocks computation, forcing devices to wait for activations to be transmitted. Displaced parallelism mitigates this bottleneck by leveraging activation similarity across steps: each step sends activations immediately, which are then used in the next step (though outdated). This “send now, receive next” strategy prevents blocking while still providing a close approximation of current data due to similarity (also termed redundancy). As visualized in Figure 2(b), experts of layer $N$ at step $T+1$ directly use activations transmitted from step $T+2$, enabling computation-communication overlap.
@@
Based on these findings, we propose DICE (Diffusion Inference with staleness-CEntric optimizations), which enhances parallel inference at the step, layer, and token granularities. DICE incorporates a three-fold approach to manage staleness effectively: (1) Interweaved Parallelism reduces step-level staleness from 2-steps to 1-step compared to displaced expert parallelism, halving the required buffer size, and achieving this without additional overhead. Detailed in Figure 2(c). (2) Selective Synchronization synchronizes only the layers most sensitive to staleness, ensuring that critical information remains up-to-date; and (3) Conditional Communication, a fine-grain strategy that adjusts communication frequency based on token importance. Together, these optimizations mitigate staleness issues while enhancing memory usage and inference efficiency, achieving up to a 1.26×1.26$\times$1.26 × speedup with minimal quality impact.
@@
  • We are the first to identify the issue of staleness in MoE-based diffusion model inference, highlighting its impact on performance.
@@
  • Based on these insights, we propose DICE, which optimizes staleness at the step, layer, and token granularities through interweaved parallelism, selective synchronization, and conditional communication, respectively. DICE outperforms state-of-the-art methods, providing a cohesive framework to improve the quality, efficiency, and scalability of parallel inference. Our code is now publicly available.
@@ 2.1 MoE-based Diffusion Models
Mixture-of-Experts (MoE) [32] is an architecture that effectively scales model capacity by partitioning the network into multiple experts, with only a subset activated per input; further detailed in supplement Section 7.1. DiT-MoE [5] integrates MoE into Diffusion Transformers to effectively scale diffusion models to 16.5 billion parameters. It also employs shared experts to capture common knowledge [29]. The largest DiT-MoE model features 32 experts and achieves state-of-the-art performance on ImageNet [2].
@@ 2.2 Displaced Parallelism
Displaced Parallelism, introduced by DistriFusion [20], leverages activation similarity between successive diffusion steps to asynchronously transfer activations, overlapping communication and computation for speedup. Activations computed in the current step are sent asynchronously and used in the next step; expanded upon in supplement Section 7.2. However, this approach introduces staleness due to deferred communication, leading to quality degradation, especially in expert parallelism. To address this, we introduce staleness-centric optimizations that enhance efficiency and maintain output quality. PipeFusion [36] and AsyncDiff [1] also exploit activation similarity to accelerate diffusion inference. In addition to displaced parallelism, caching methods have been proposed as alternative optimizations targeting U-Net [25, 37, 7, 33], DiT [31], and learned caching [24].
@@ 2.3 Expert Parallelism
Expert Parallelism [16] scales MoE models by distributing expert weights across devices while replicating non-expert layers. The input batch is split among devices (similar to data parallelism), and tokens (patches) are routed via all-to-all communication, enabling large-scale deployment.
Figure 3: Visualization of expert parallelism. Each device holds a subset of experts (FFNs) and processes a portion of the input batch. Different colors denote distinct data samples.
MoE’s strong performance and reliance on expert parallelism have motivated optimizations. FasterMoE [9] employs pipelined all-to-all communication and expert shadowing, DeepSpeed-MoE [29] leverages efficient communication and model compression, and BASE Layers [17] use distributed linear assignment for balanced token-expert allocation; additional work addresses communication via topology-aware token routing [9, 19].
However, these methods are not tailored for diffusion models and do not tackle the unique challenges in diffusion processes. Our work fills this gap by introducing staleness-centric optimizations for MoE-based diffusion models, achieving improved quality and efficiency.
@@ 3 Motivation
Our motivation is based on three key observations: first, the communication bottleneck in expert parallelism necessitate optimization; second, we observed high redundancy in input/output and routing activation, suggesting the feasibility of optimization through asynchronous communication; and third, we identified staleness as the core issue, presenting the main challenge to performance.
Communication bottleneck in expert parallelism. All-to-all operations in expert parallelism introduce significant communication overhead, presenting a major bottleneck [9, 17, 39, 18]. Our evaluation of DiT-MoE-XL on 8 GPUs shows that all-to-all communication accounts for a substantial portion of the total inference time. Specifically, for batch sizes 4, 8, and 16, the all-to-all communication times were 15.91 seconds (61.7% of the total time), 28.99 seconds (69.8%), and 54.94 seconds (73.3%), respectively, highlighting the necessity to mitigate the communication inefficiency.
Figure 4: Step-wise similarity heatmaps in DiT-MoE, where both axes represent diffusion steps. We use a one-hot encoding to represent routing assignments.
Redundancy in MoE-based Diffusion Models. We identify significant computational redundancy in DiT-MoE’s expert routing: adjacent diffusion steps exhibit similar token-expert assignments (Figure LABEL:fig:activation_similarity) and activation value. This strong consistency in both activation  [25, 24] and routing decisions persists across denoising steps. Such redundancy enables asynchronous communication strategies in expert-parallel systems.
Staleness degrades performance. We discovered that using outdated (staled) activations due to communication delays in expert parallelism severely degrades model performance. Displaced parallelism induces a two-step staleness in expert parallelism, which causes a notable drop in image quality: the FID score increases from 5.31 to 8.27(Sec. 5). Our staleness-centric optimizations decrease staleness to just one step, along with efficient staleness trade-offs, leading to significant speedup and quality improvement.
@@ 4 Methodology
Figure 5: Overview of DICE. The blue path highlights layer N’s dataflow. Interweaved Parallelism interleaves operations; This execution pattern reduces staleness from two steps to one, half of displaced expert parallelism [20]. Selective Synchronization targets staleness-vulnerable deeper layers, and Conditional Communication prioritizes important tokens based on router scores.
Efficient parallel inference for MoE-based diffusion models must address staleness—the use of outdated activations caused by asynchronous communication. While displaced expert parallelism [20] accelerates inference through overlapping, its asynchronous nature induces severe staleness, compromising output quality.
@@ 4.1 Interweaved Parallelism
State-of-the-art methods use displaced parallelism [20, 36] to overlap communication and computation in distributed diffusion inference. By asynchronously initiating communication in one step and using the results in the subsequent step, it reduces blocking time.
@@
[…]
@@
[…]
@@
We emphasize that interweaved parallelism is a free-lunch optimization over the state-of-the-art method, halving staleness for better image quality, reducing buffer size, and incurring no additional latency.
@@
Figure 7: Illustration of Conditional Communication. High-scoring tokens are transmitted every step to maintain freshness, while lower-scoring tokens reuse cached value and update every few steps (stride=2 in this example).“Unselected” refers to scores outside the top-K choices(K=activated experts per token), meaning the corresponding experts are not activated and no data should be transmitted.
@@ 4.2 Selectively Synchronize Vulnerable Layers
At the layer level, selectively synchronizing vulnerable layers can significantly improve image quality. Our analysis reveals that the impact of staleness exhibits layer-wise asymmetry in MoE-based diffusion models: shallow experts extracting low-level features are inherently robust to asynchronous communication, while experts in deeper layers handling high-level semantics prove fundamentally vulnerable to activation staleness. This architectural property aligns with DeepSpeed-MoE’s observation in language models [29] where deeper layers benefit more from MoE.
Guided by this understanding, we propose Selective Synchronization, which synchronizes only the more vulnerable deeper layers while allowing shallow layers to continue asynchronously. This targeted approach ensures that deeper layers benefit from up-to-date information without the need for full synchronization. Experimental validation in Figure 6 confirms the layer-wise vulnerability pattern. Ablation studies (Sec. 5.4) demonstrating how our partial synchronization achieves optimal quality-efficiency tradeoffs.
@@ 4.3 Freshness-Latency Trade-Off via Conditional Communication
To quantify staleness impact, For token $i$ assigned to expert $e$ with expert activations $\mathbf{h}_{i}^{e}$ and router scores $s_{i}^{e}$, the output $\mathbf{y}_{i}=\sum_{e=1}^{E}s_{i}^{e}\mathbf{h}_{i}^{e}$ with L2-norm $\|\mathbf{y}_{i}\|$. The gradient:
@@
Our training-free solution maintains freshness for critical tokens while tolerating staleness in less impactful ones, as illustrated in Figure 7; elaborated further in supplement Section 9. It achieves substantial communication reduction with minimal quality trade-off, as validated in Section 5.4.
@@ 5.1 Setups
We provide additional results on different hardware in supplement Section 5.
Models and Datasets. We evaluate our approach on two configurations of the DiT-MoE model [6]: DiT-MoE-XL, which is configured with 8 experts across 28 layers, and DiT-MoE-G, with 16 experts across 40 layers, both containing 2 extra shared experts. DiT-MoE is trained on ImageNet, which we use as the benchmark for evaluation.
Baselines. To evaluate the effectiveness of our approach, we compare it against three baselines: DistriFusion [20] (displaced sequence parallelism), a state-of-the-art method for distributed diffusion models employing patch parallelism; Expert Parallelism [16], the standard approach for parallelizing MoE models; and displaced expert parallelism, the displaced variant [20] of expert parallelism, which overlaps communication and computation.
Our main results use a 50-step configuration, which is hardcoded in the DiT-MoE codebase, for fair comparisons. We also provide additional results for 10—and 20-step settings, showing that DICE is even more effective under fewer steps.
Metrics. We assess the performance of our method using several well-established metrics, including Fréchet Inception Distance (FID) [10], Sliced Fréchet Inception Distance (sFID) [26], Inception Score (IS) [30], as well as Precision and Recall [15]. FID and sFID measures distributional alignment, while IS, precision, and recall assess various aspects of sample quality.
Implementation Details. Our implementation is based on PyTorch 2.0.0+cu118 [27]. We utilize Rectified Flow [22, 11] for the DiT-MoE-XL and G models, as they are specifically trained to support this approach. The core codebase builds upon the original DiT-MoE implementation [4]. For expert parallelism, we referenced FastMoE [8]. Sequence Parallelism [21] and DistriFusion [20] are adapted to DiT-MoE. We provide more details in supplement Section 9.
Class-Conditional ImageNet 256$\times$256

### v4

@@ Abstract
Mixture-of-Experts-based (MoE-based) diffusion models demonstrate remarkable scalability in high-fidelity image generation, yet their reliance on expert parallelism introduces critical communication bottlenecks. State-of-the-art methods alleviate such overhead in parallel diffusion inference through computation-communication overlapping, termed displaced parallelism. However, we identify that these techniques induce severe staleness — the usage of outdated activations from previous timesteps that significantly degrades quality, especially in expert-parallel scenarios. We tackle this fundamental tension and propose DICE, a staleness-centric optimization framework with a three-fold approach: (1) Interweaved Parallelism introduces staggered pipelines, effectively halving step-level staleness for free; (2) Selective Synchronization operates at layer-level and protects layers vulnerable to staled activations; and (3) Conditional Communication, a token-level, training-free method that dynamically adjusts communication frequency based on token importance. Together, these strategies effectively reduce staleness, achieving $1.26\times$ speedup with minimal quality degradation. Empirical results establish DICE as an effective and scalable solution. Our code is available at https://github.com/Cobalt-27/DICE
Figure 1: We introduce DICE, a framework designed to enhance scalability and efficiency in MoE-based diffusion models. Existing asynchronous methods accelerate parallel inference but suffer from outdated activations (staleness). DICE tackles this issue with step-, layer-, and token-level optimizations, reducing staleness and achieving up to $1.26\times$ speedup with minimal quality loss. Left: Architectural comparison; Right: Visual quality results.
[…]
@@ 1 Introduction
Figure 2: Execution flows of (a) synchronous expert parallelism (no staleness); (b) displaced variant of expert parallelism (two-step staleness, proposed in DistriFusion) [distri] and (c) our interweaved parallelism (one-step staleness). Each color denotes a different layer, showing how interweaving reduces staleness by staggering operations within the same step. The blue path highlights layer NN, which spans two more steps in displaced parallelism but only one in interweaved. FID and speedup refer to interweaved parallelism alone (Section 5.5).
Diffusion models [ddpm, dhariwal2021diffusion] have revolutionized the field of generative modeling, enabling the creation of high-fidelity images [restoration, score] and videos [video, luo2023videofusion], rivaling human creativity in quality and detail. Among the architectures, the Diffusion Transformer (DiT) [dit, transformer, qdit] with Mixture-of-Experts (MoE) [moe] stands out due to its effectiveness in scalability. MoE enables large diffusion models to scale up with sub-linear cost increases. It divides the model into multiple experts, with only a subset activated for each input. This selective activation reduces computation, achieving high performance at a lower cost. As demonstrated by DiT-MoE [ditmoe, rap], MoE-based diffusion models have been scaled to 16 billion parameters with superior quality.
Scaling MoE-based diffusion models necessitates expert parallelism [gshard, deepspeedmoe] to handle their extensive memory footprint; however, it also introduces significant communication overhead. By distributing expert weights across multiple GPUs, expert parallelism substantially reduces memory usage but incurs significant communication overhead [fastermoe, baselayers, interlayer, lina]. As outlined in Figure 2(a), two all-to-all communications (dispatch + combine) are required per layer. This bottleneck presents a major challenge for scaling diffusion models.
To address communication challenges in parallel inference, state-of-the-art methods employ Displaced Parallelism [distri]. In standard synchronous execution, communication blocks computation, forcing devices to wait for activations to be transmitted. Displaced parallelism mitigates this bottleneck by leveraging activation similarity across steps: each step sends activations immediately, which are then used in the next step (though outdated). This “send now, receive next” strategy prevents blocking while still providing a close approximation of current data due to similarity (also termed redundancy). As visualized in Figure 2(b), experts of layer $N$ at step $T+1$ directly use activations transmitted from step $T+2$, enabling computation-communication overlap.
@@
Based on these findings, we propose DICE (Diffusion Inference with staleness-CEntric optimizations), which enhances parallel inference at the step, layer, and token granularities. DICE incorporates a three-fold approach to manage staleness effectively: (1) Interweaved Parallelism reduces step-level staleness from 2-steps to 1-step compared to displaced expert parallelism, halving the required buffer size, and achieving this without additional overhead. Detailed in Figure 2(c). (2) Selective Synchronization synchronizes only the layers most sensitive to staleness, ensuring that critical information remains up-to-date; and (3) Conditional Communication, a fine-grain strategy that adjusts communication frequency based on token importance. Together, these optimizations mitigate staleness issues while enhancing memory usage and inference efficiency, achieving up to a $1.26\times$ speedup with minimal quality impact.
@@
  • We identify the issue of staleness in MoE-based diffusion model inference, highlighting its impact on performance.
@@
  • Based on these insights, we propose DICE, which optimizes staleness at the step, layer, and token granularities through interweaved parallelism, selective synchronization, and conditional communication, respectively. DICE outperforms state-of-the-art methods, providing a cohesive framework to improve the quality, efficiency, and scalability of parallel inference. Our code is publicly available.
@@ 2.1 MoE-based Diffusion Models
Mixture-of-Experts (MoE) [moe] is an architecture that effectively scales model capacity by partitioning the network into multiple experts, with only a subset activated per input; further detailed in supplement Section 7.1. DiT-MoE [ditmoe] integrates MoE into Diffusion Transformers to effectively scale diffusion models to 16.5 billion parameters. It also employs shared experts to capture common knowledge [deepspeedmoe]. The largest DiT-MoE model features 32 experts and achieves state-of-the-art performance on ImageNet [imagenet].
@@ 2.2 Displaced Parallelism
Displaced Parallelism, introduced by DistriFusion [distri], leverages activation similarity between successive diffusion steps to asynchronously transfer activations, overlapping communication and computation for speedup. Activations computed in the current step are sent asynchronously and used in the next step. This mechanism is expanded upon in supplement Section  7.2. However, this approach introduces staleness due to deferred communication, leading to quality degradation, especially in expert parallelism. To address this, we introduce staleness-centric optimizations that enhance efficiency and maintain output quality. PipeFusion [pipefusion] and AsyncDiff [asyncdiff] also exploit activation similarity to accelerate diffusion inference. In addition to displaced parallelism, caching is an alternative optimization targeting both U-Net [deepcache, wimbauer2024cache, habibian2024clockwork, so2023frdiff] and DiT [fora, learn2cache].
@@ 2.3 Expert Parallelism
Expert Parallelism [gshard] scales MoE models by distributing expert weights across devices while replicating non-expert layers. The input batch is split among devices (similar to data parallelism), and tokens (patches) are routed via all-to-all communication, enabling large-scale deployment. Compared to other parallelism strategies[megatron, lu2023dagc, lu2024tmc, dist_survey], expert parallelism is inherently more compatible with MoE for its alignment with the architecture.
Figure 3: Visualization of expert parallelism. Each device holds a subset of experts (Feed-Forward Networks, FFNs) and processes a portion of the input batch. Different colors denote distinct data samples.
MoE’s strong performance and reliance on expert parallelism have motivated optimizations. FasterMoE [fastermoe] employs pipelined all-to-all communication and expert shadowing, DeepSpeed-MoE [deepspeedmoe] leverages efficient communication and model compression; additional work addresses communication via topology-aware token routing [locmoe, fastermoe].
However, these methods are not tailored for diffusion models and do not tackle the unique challenges in diffusion processes. Our work fills this gap by introducing staleness-centric optimizations.
@@ 3 Motivation
Our motivation is based on three key observations: first, expert parallelism suffers from a critical communication bottleneck that necessitates optimization; second, MoE-based diffusion models exhibit high redundancy in their activations and routing, suggesting the feasibility of asynchronous communication; and third, we identify staleness, a direct consequence of such asynchrony, as the core challenge to performance.
Communication bottleneck in expert parallelism. All-to-all operations in expert parallelism introduce significant communication overhead, presenting a major bottleneck [fastermoe, baselayers, interlayer, lina]. Our evaluation of DiT-MoE-XL on 8 GPUs shows that all-to-all communication accounts for a substantial portion of the total inference time. Specifically, for batch sizes 4, 8, and 16, the all-to-all communication times were 15.91 seconds (61.7% of the total time), 28.99 seconds (69.8%), and 54.94 seconds (73.3%), respectively, highlighting the necessity to mitigate the communication inefficiency.
Figure 4: Step-wise similarity heatmaps (cosine similarity between diffusion steps) in DiT-MoE, with both axes as diffusion steps. Results are shown for Layer 5 and 25, using one-hot routing assignments for similarity computation.
Redundancy in MoE-based Diffusion Models. We identify significant computational redundancy in DiT-MoE’s expert routing: adjacent diffusion steps exhibit similar token-expert assignments (Figure 4) and activation values. This strong consistency in both activation  [deepcache, learn2cache] and routing decisions persists across denoising steps. Such redundancy enables asynchronous communication strategies in expert-parallel systems.
Staleness degrades performance. We discovered that using outdated (stale) activations due to communication delays in expert parallelism severely degrades model performance. Displaced parallelism induces a two-step staleness in expert parallelism, which causes a drop in image quality: the FID score increases from 5.31 to 8.27 (Section 5).
@@ 4 Methodology
Figure 5: Overview of DICE. The blue path highlights layer N’s dataflow. Interweaved Parallelism interleaves operations; This execution pattern reducing the steps that cause staleness from two to one, half of displaced expert parallelism [distri]. Selective Synchronization targets staleness-vulnerable deeper layers, and Conditional Communication prioritizes important tokens based on router scores.
Efficient parallel inference for MoE-based diffusion models must address staleness—the use of outdated activations caused by asynchronous communication. While displaced expert parallelism [distri] accelerates inference through overlapping, its asynchronous nature induces severe staleness, compromising output quality.
@@ 4.1 Interweaved Parallelism
State-of-the-art methods use displaced parallelism [distri, pipefusion] to overlap communication and computation in distributed diffusion inference. By asynchronously initiating communication in one step and using the results in the subsequent step, it reduces blocking time.
@@
$\text{Staleness}_{\,\ \text{displaced}}=\underbrace{1}_{\text{dispatch}}+\underbrace{1}_{\text{combine}}=2\text{-steps}.$
@@
$\text{Staleness}_{\,\ \text{interweaved}}=\underbrace{0}_{\text{dispatch}}+\underbrace{1}_{\text{combine}}=1\text{-step}.$
@@
Interweaved parallelism serves as a free-lunch optimization over displaced parallelism, halving staleness for better image quality, reducing buffer size, while achieving the same degree of overlap.
@@
Figure 7: Illustration of Conditional Communication. High-scoring tokens are transmitted every step to maintain freshness, while lower-scoring tokens reuse cached value and update every few steps (stride=2 in this example).”Unselected” refers to scores outside the top-K choices(K is the number of activated experts per token), indicating that the corresponding experts are not activated and no data should be transmitted.
@@ 4.2 Selectively Synchronize Vulnerable Layers
At the layer level, selectively synchronizing vulnerable layers can significantly improve image quality. Our analysis reveals that the impact of staleness exhibits layer-wise asymmetry in MoE-based diffusion models: shallow experts extracting low-level features are inherently robust to asynchronous communication, while experts in deeper layers handling high-level semantics prove fundamentally vulnerable to activation staleness. This architectural property aligns with DeepSpeed-MoE’s observation in language models [deepspeedmoe] where deeper layers benefit more from MoE.
Guided by this understanding, we propose Selective Synchronization, which synchronizes only the more vulnerable deeper layers while allowing shallow layers to continue asynchronously. This targeted approach ensures that deeper layers benefit from up-to-date information without the need for full synchronization. Experimental validation in Figure 6 confirms the layer-wise vulnerability pattern. Ablation studies (Section 5.4) demonstrate how our partial synchronization achieves optimal quality-efficiency tradeoffs.
@@ 4.3 Freshness-Latency Trade-Off via Conditional Communication
To quantify staleness impact, consider a token $i$ assigned to expert $e$ with expert activations $\mathbf{h}_{i}^{e}$ and router scores $s_{i}^{e}$, the output $\mathbf{y}_{i}=\sum_{e=1}^{E}s_{i}^{e}\mathbf{h}_{i}^{e}$ with L2-norm $\|\mathbf{y}_{i}\|$. The gradient:
@@
Our training-free solution maintains freshness for critical tokens while tolerating staleness in less impactful ones, as shown in Figure 7; elaborated in supplement Section 9. It achieves substantial communication reduction with minimal quality trade-off, as validated in Section 5.4.
@@ 5.1 Setups
Models and Datasets. We evaluate our approach on two configurations of the DiT-MoE model [ditmoe-hf]: DiT-MoE-XL, which is configured with 8 experts across 28 layers, and DiT-MoE-G, with 16 experts across 40 layers, both containing 2 extra shared experts. DiT-MoE is trained on ImageNet, which we use as the benchmark for evaluation.
Baselines. To evaluate the effectiveness of our approach, we compare it against three baselines: DistriFusion [distri] (displaced sequence parallelism), a state-of-the-art method for distributed diffusion models employing patch parallelism; Expert Parallelism [gshard], the standard approach for parallelizing MoE models; and displaced expert parallelism, the displaced variant [distri] of expert parallelism, which overlaps communication and computation.
Our main results use a 50-step configuration, which is hardcoded in the DiT-MoE codebase, for fair comparisons. We also provide additional results for 10-and 20-step settings, showing that DICE is even more effective under fewer steps.
Metrics. We assess the performance of our method using several well-established metrics, including Fréchet Inception Distance (FID) [fid], Sliced Fréchet Inception Distance (sFID) [sfid], Inception Score (IS) [is], as well as Precision and Recall [prec-recall].
Implementation Details. Our implementation is based on PyTorch 2.0.0+cu118 [torch]. We utilize Rectified Flow [rf, cfg] for the DiT-MoE-XL and G models, as they are specifically trained to support this approach. The core codebase builds upon the original DiT-MoE implementation [ditmoe-git]. For expert parallelism, we referenced FastMoE [fastmoe]. Sequence Parallelism [seqpara] and DistriFusion [distri] are adapted to DiT-MoE. We provide more details in supplement Section 9.
Class-Conditional ImageNet $\times256$